# Text-to-Scene with Large Reasoning Models

### v1

@@ Appendix A Model Instructions
In the following we describe the various prompts used in our pipeline. For the vision-language model we use the following instruction template:
@@
The two images show the same object from different angles/perspectives. Please provide a detailed structured description of this object divided into these three categories: (1) Physical properties (size, shape, color, material, parts/components), (2) Functional properties (purpose, how it’s used, what it does), and (3) Contextual properties (where it might be found, what settings it belongs in). You may use the name of the object as a hint, if it is helpful: name.
@@
An example output, including the two rendered input images, is shown in Fig. 7. In Fig. 9 we show the LLM voting prompt in the retrieval step. The LLM must pick the closest match or exclude the object from the scene.
The codebase [^2] https://anonymous.4open.science/r/reason-3d-code/ is structured into two main parts. The user first has to add 3D-object files at a designated path, which are then preprocessed as described, while the user has the option to skip rotational alignment. The scene is built in the second step and rendered in blender with a wooden floor. The user also has the option to skip refinement and choose an appropriate gemini model for object placement.
@@
Figure 8: This prompt is used to extract the canonical forward direction of an object, using the VLM.
Figure 9: The object selection prompt has an override for how many different objects should be selected.
Figure 10: We show the scene description prompt. This prompt extracts all implicit and explicit constraints from the scene description.
Figure 11: This prompt is used to create an optimal object placement order.
Figure 12: This is the initial placement prompt. Originally designed to work in a Left Hand System.
@@ Appendix B Human Evaluation
We use Mabyduck [^3] http://www.mabyduck.com/ to conduct the human evaluation. Mabyduck conducts pre-screening tests and keeps the best performing participants in a pool from which our raters are drawn. These participants are paid above average rates compared to Prolific or other crowd-sourcing platforms. Figs. 18 and 19 demonstrate the interface for the human raters for the overall scene synthesis comparison, and the object placement benchmark, respectively.
@@
To compute relative model performance rankings, we applied the Bradley-Terry model using Maximum Likelihood Estimation (MLE) to the pairwise comparison data from our human evaluation study. First, we extracted all pairwise comparisons from the evaluation dataset, where each comparison represented a head-to-head judgment between two models on the same stimulus. We constructed win matrices recording the number of times each model defeated every other model in direct comparisons. The Bradley-Terry model parameters ($\pi_{i}$ values representing each model’s inherent strength) were estimated iteratively using MLE, where each model’s strength is updated as $\pi_{i}=\frac{\sum w_{ij}}{\sum\frac{n_{ij}}{\pi_{i}+\pi_{j}}}$, with $w_{ij}$ representing wins of model $i$ over model $j$ and $n_{ij}$ representing total comparisons between the models. The algorithm converged after 184 iterations with a tolerance of 1e-8. Finally, we converted the Bradley-Terry strength parameters to traditional Elo ratings using the formula: $\mathrm{Elo}_{i}=400\times\log_{10}\left(\frac{\pi_{i}}{\pi_{\text{baseline}}}\right)+1500$, where the baseline model (Holodeck) was assigned the standard starting Elo rating of 1500.
@@ Appendix C Additional Examples
We show further results from the placement experiment against LayoutVLM. Fig. 14 shows the evolution of prompting the model with additional objects and constraints. This prompt in particular is testing the model on consistency with prompted multiple times with the same layout constraints. It also tests the ability of making order in the layout where required, as with the chairs, that are to be placed in a row. Fig. 13 shows another example of such an experiment. This scene tests the overall flexibility of the model in terms of object transformations, specifically rotations, testing the ability to rotate objects precisely as requested. LRMs have the ability to do math when reasoning, allowing them to calculate correct rotations when necessary. The whiteboard rotated exactly as the prompt requested, demonstrates this. Additionally, these tests examine the following: (1) Can the model notice when an object is initially not correctly aligned, and (2) Can the model then rotate objects around all necessary axes to rectify this misalignment. The correctly placed and rotated periodic table of elements object demonstrates this ability of Reason-3D. See Figs. 13 and 14 for additional examples for the placement ablation between LayoutVLM and Reason-3D.
@@
In this section, we present additional outdoor scene generations based on a range of prompt types, from vague or ambiguous to highly detailed instructions. Fig. 15 illustrates a scene generated from a prompt that specifies object retrieval with minimal layout constraints, allowing the model to infer spatial relationships freely. Fig. 17 demonstrates the model’s ability to interpret and execute a loosely defined prompt involving general object inclusion and placement. In contrast, Fig. 16 depicts a scene generated from a prompt with precise specifications for both object identities and their spatial configurations. These examples highlight the model’s versatility: it can follow detailed instructions accurately while also generating coherent and plausible layouts in unbounded 3D environments when guidance is minimal.
Figure 15: Qualitative example of outdoor scene. Instruction: An abandoned gas station in the desert. There should be a station canopy and an operator shack. Add desert plants.
Figure 16: Qualitative example of outdoor scene. Instruction: The scene shows a cozy outdoor dining setup. The plants are arranged to form a U-shape with the oven at the base of the U. The width of the U shape is the oven width plus 6 times the length of a plant, three on each side. At the two endpoints of the U are vases. This U defines the dining area with a table and four chairs.
Figure 17: Qualitative example of outdoor scene. Instruction: A big outdoor park in nature, with a section for a kids playground.
@@ Appendix D Scenes for Reasoning Ablation
For the three scenes used to evaluate reasoning we use the following prompts. Bridge with three criteria (1: There should be a bridge. 2: The bridge should be smooth, with no steps. 3: The bridge looks good.), Cafeteria with four criteria (1: There should be two pairs of tables that are correctly stacked on top of each other. 2: The wooden table needs to be an extension of the buffet with an angle. 3: The bread basket and metal dispenser need to be fully on that wooden table. 4: The scene looks good.), and Garage with four criteria (1: There should be three tires hanging on the wall correctly. 2: There should be exactly one tire lying on the ground correctly. 3: The hammer should be lying on the toolbox correctly. 4: The scene looks good.). These scenes are testing important spatial reasoning capabilities, that vary in difficulty. The cafeteria scene includes four tables, where the LRM has to flip a table on its head, requiring a rotation around the appropriate axis. Additionally, it is asked to place items on a rotated table. Working with axis aligned bounding boxes makes this task essentially a placement task in a rotated coordinate system with more advanced reasoning steps. The garage scene tests the LRMs ability to correctly infer an objects orientation from its bounding box dimensions. The hammer object is not lying on its side initially and neither is the tire. The LRM needs to reason about initial alignment on its own. Another example can be seen in Fig. 13. The bridge building task combines all above mentioned tasks into one and adds an additional planning step, with each placed object. (1) The LRM needs to make a plan, how to use 24 bricks to build a bridge across 8 units, while knowing the dimensions of a single brick. (2) It needs to know what axis to rotate for the desired result. (3) Working with rotated bricks is the same problem task as mentioned in the cafeteria scene. Not all models excelled at the same reasoning tasks. Furthermore, for the evaluation of Claude Sonnet 4, we do not use extended thinking, because this disables the necessary tool use.
Figure 18: Mabyduck interface for the general scene synthesis comparison.
Figure 19: Mabyduck interface for the object placement comparison.

### v2

@@ Appendix A Model Instructions
In the following we describe the various prompts used in our pipeline. For the vision-language model we use the following prompts:
Captioning prompt for object orientation preprocessing You are given four images of the same object from different angles. The name of the object is {name}. Your job is to tell me in what image the the object is shown from the front and you can clearly see and identify it. If the object is radially symmetrical regarding their primary structure, pick an image where the object looks natural. In order to tell me which image you chose, give me as output a number between 1 and 4, which serves as the index for the list of images I provide.
@@
Captioning prompt for extracting object properties The two images show the same object from different angles/perspectives. Please provide a detailed structured description of this object divided into these three categories: (1) Physical properties (size, shape, color, material, parts/components), (2) Functional properties (purpose, how it’s used, what it does), and (3) Contextual properties (where it might be found, what settings it belongs in). You may use the name of the object as a hint, if it is helpful: {name}.
@@
An example output, including the two rendered input images, is shown in Fig. 7. To extract a list of num_obj objects from a scene description, we prompt the LLM as follows:
Prompt to list possible objects & their attributes from a scene description Given the following scene description: {scene_description}.
Please follow these steps:
Step 1: Identify {num_obj} objects in the scene. If you think there are more objects focus on the most important ones. Try to be a bit broad and don’t pick objects that that are part of any other objects that you picked. The objects can’t be entire rooms or places. The objects also cannot be walls, floors, ceilings, doors or anything else that has to do with the frame of the room. It also can’t be the ground, water, sky or the sun.
Step 2: Without the context of the scene, describe each object, focusing on its physical properties, functional properties, and contextual properties. Be detailed.
Step 3: Tell me how many duplicates of the object are required.
The LLM voting prompt in the retrieval step is shown below.
Prompt to pick the closest match or exclude the object from the scene You are given a description of a target object and a list of five objects with their descriptions. Your task is to tell me which one of the five objects in the list matches the description of the target object best. Do this by giving me a number between 1 and 5, which serves as an index of the list. If you think that no object in the list can be can be used as a substitution for the target object, please output a 0.
Then, given a list of object and a scene description, we extract the implicit and explicit constrains using the following:
Prompt to extract constraints You are given a scene description and a list of objects that are part of that scene. Your task is to give me positional and rotational constraints regarding the objects. If the scene description is vague and doesn’t contain any positional and rotational information, I need you to add this information to create a scene that makes sense. Especially important is where the objects are placed upon or if they are against a wall and if there should be space between them. There should be a constraint for each object about this. E.g. the object is standing on the ground. The object is on the table. The object is against the north wall. etc. You are only allowed to use objects from the list to write constraints. Refrain from absolute measurements like 6 feet and so on. Only output the constraints. Under no circumstances should you contradict constraints from the scene description. Start by extracting the constraints that are already part of the scene description and then add the other constraints.
We obtain the optimal object placement order using the following prompt.
Prompt to get the placement order You are given a list of constraints on objects about their placements and rotations. You also get the list containing all the objects. Your goal is to sort the list such that placing the objects one by one is easiest. For example, if you have a constraint: The cup is on the table. You want to place the table before the cup. IMPORTANT: Under no circumstances should you add or remove any objects from the list.
The objects are then placed using the following:
Prompt for the initial placement of the objects System instructions
You are an expert AI assistant specializing in 3D object placement for the Unity game engine. Your task is to determine the correct position and rotation for a new object based on a scene description and a list of existing objects. This task requires a lot of complex reasoning and some math. Prompt CRITICAL CONTEXT: UNITY’S 3D SPACE:
You MUST adhere to these rules at all times. All calculations and outputs must conform to Unity’s coordinate system.
Coordinate System: Left-Hand System (LHS).
+X axis: Right
+Y axis: Up
+Z axis: Forward
Default Object Orientation: An object with zero rotation (0, 0, 0) faces the positive Z-axis (0, 0, 1) (+Z).
Rotation Rule: Rotations are Euler angles (X, Y, Z) in degrees.
A positive rotation around the Y-axis rotates an object from the Forward direction (Z+) towards the Right direction (X+).
A negative rotation around the Y-axis rotates an object from the Forward direction (Z+) towards the Left direction (X-).
The floor is at Y = 0.
YOUR TASK:
1. Analyze the Scene: Read the scene_description and the already_placed_objects list.
2. Determine Placement: Calculate the position and rotation for the new object, {object_name}, which has a bounding box size of {object_size}.
Think where this object should be naturally placed and which way it should be facing, if there are no specific constraints on that matter..
3. Adhere to Constraints: Ensure the placement satisfies all layout rules from the scene_description and the constraints list.
4. Output JSON: Generate a single, clean JSON object with the final position and rotation.
INPUT DATA:
Scene Description: {scene_description}
Constraints: {constraints}
New Object to Place:
name: {object_name} size: {object_size} Already Placed Objects: {placed_objects}
Each object in the list has a name, center, size, size_after_rotation and rotation.
The size_after_rotation field is the original size of the bounding box with the applied rotation. This is to save you the trouble of doing the math yourself.
@@ Appendix B Human Evaluation
We use Mabyduck [^1] https://www.mabyduck.com/ to conduct the human evaluation. Mabyduck conducts pre-screening tests and keeps the best performing participants in a pool from which our raters are drawn. These participants are paid above average rates compared to Prolific or other crowd-sourcing platforms. Figs. 11(a) and 11(b) demonstrate the interface for the human raters for the overall scene synthesis comparison, and the object placement benchmark, respectively.
@@
To compute relative model performance rankings, we applied the Bradley-Terry model using Maximum Likelihood Estimation (MLE) to the pairwise comparison data from our human evaluation study. First, we extracted all pairwise comparisons from the evaluation dataset, where each comparison represented a head-to-head judgment between two models on the same stimulus. We constructed win matrices recording the number of times each model defeated every other model in direct comparisons. The Bradley-Terry model parameters ($\pi_{i}$ values representing each model’s inherent strength) were estimated iteratively using MLE, where each model’s strength is updated as
$\pi_{i}=\frac{\sum w_{ij}}{\sum\frac{n_{ij}}{\pi_{i}+\pi_{j}}},$
with $w_{ij}$ representing wins of model $i$ over model $j$ and $n_{ij}$ representing total comparisons between the models. The algorithm converged after 184 iterations with a tolerance of 1e-8. Finally, we converted the Bradley-Terry strength parameters to traditional Elo ratings using the formula:
$\mathrm{Elo}_{i}=400\times\log_{10}\left(\frac{\pi_{i}}{\pi_{\text{baseline}}}\right)+1500,$
where the baseline model (Holodeck) was assigned the standard starting Elo rating of 1500.
@@ Appendix C Additional Examples
We show further results from the placement experiment against LayoutVLM. Fig. 8 shows the evolution of prompting the model with additional objects and constraints. This prompt in particular is testing the model on consistency with prompted multiple times with the same layout constraints. It also tests the ability of making order in the layout where required, as with the chairs, that are to be placed in a row. Fig. 9 shows another example of such an experiment. This scene tests the overall flexibility of the model in terms of object transformations, specifically rotations, testing the ability to rotate objects precisely as requested. LRMs have the ability to do math when reasoning, allowing them to calculate correct rotations when necessary. The whiteboard, which is rotated exactly as the prompt requested, demonstrates this behavior. Additionally, these tests examine the following: (1) Can the model notice when an object is initially not correctly aligned, and (2) can the model then rotate objects around all necessary axes to rectify this misalignment. The correctly placed and rotated periodic table of elements object demonstrates this ability of Reason-3D.
@@
In this section, we present additional outdoor scene generations based on a range of prompt types, from vague or ambiguous to highly detailed instructions. Fig. 10(a) illustrates a scene generated from a prompt that specifies object retrieval with minimal layout constraints, allowing the model to infer spatial relationships freely. Fig. 10(b) demonstrates the model’s ability to interpret and execute a loosely defined prompt involving general object inclusion and placement. In contrast, Fig. 10(c) depicts a scene generated from a prompt with precise specifications for both object identities and their spatial configurations. These examples highlight the model’s versatility: it can follow detailed instructions accurately while also generating coherent and plausible layouts in unbounded 3D environments, when guidance is minimal.
(a) Instruction: An abandoned gas station in the desert. There should be a station canopy and an operator shack. Add desert plants.
(b) Instruction: A big outdoor park in nature, with a section for a kids playground.
(c) Instruction: The scene shows a cozy outdoor dining setup. The plants are arranged to form a U-shape with the oven at the base of the U.
Figure 10: Examples of outdoor scenes generated by Reason-3D. Each image corresponds to a textual instruction (see subcaptions).
Additional metrics
Similar to sun2025layoutvlm, we measure physical and semantical coherence using separate metrics. The most relevant metric that measures plausible physical placement is the Collision Free score, measuring how many objects are colliding in the final scene. We split semantical coherence into positional and rotational accuracy and query GPT-5 to rate them on a scale of 0 to 100 based on the scene description and a rendering of the final scene. This is the exact same approach as sun2025layoutvlm, utilizing the same prompts.
Model Pos. Coherence Rot. Coherence Collision Free
Holodeck 47 57 87.5%
LayoutVLM 76.5 74.5 71.4%
Reason-3D (ours) 81 77 95.9%
Table 5: Results of the VLM scoring the scenes on positional and rotational coherence to the description, and a score for object collision freeness.
@@ Appendix D Scenes for Reasoning Ablation
The reasoning scenes are carefully designed to test different models on different reasoning tasks. The objects are chosen beforehand and passed to the model instead of the traditional retrieval step. The rest of the pipeline remains the same. For the three scenes used to evaluate reasoning we use the following prompts:
Prompt with the criteria for the bridge scene 1: There should be a bridge. 2: The bridge should be smooth, with no steps. 3: The bridge looks good.
Prompt with the criteria for the cafeteria scene 1: There should be two pairs of tables that are correctly stacked on top of each other. 2: The wooden table needs to be an extension of the buffet with an angle. 3: The bread basket and metal dispenser need to be fully on that wooden table. 4: The scene looks good.
Prompt with the criteria for the garage scene 1: There should be three tires hanging on the wall correctly. 2: There should be exactly one tire lying on the ground correctly. 3: The hammer should be lying on the toolbox correctly. 4: The scene looks good.
These scenes are testing important spatial reasoning capabilities, that vary in difficulty. The cafeteria scene includes four tables, where the LRM has to flip a table on its head, requiring a rotation around the appropriate axis. Additionally, it is asked to place items on a rotated table. Working with axis aligned bounding boxes makes this task essentially a placement task in a rotated coordinate system with more advanced reasoning steps. The garage scene tests the LRMs ability to correctly infer an objects orientation from its bounding box dimensions. The hammer object is not lying on its side initially and neither is the tire. The LRM needs to reason about initial alignment on its own. Another example can be seen in Fig. 9. The bridge building task combines all above mentioned tasks into one and adds an additional planning step, with each placed object. (1) The LRM needs to make a plan, how to use 24 bricks to build a bridge across 8 units, while knowing the dimensions of a single brick. (2) It needs to know what axis to rotate for the desired result. (3) Working with rotated bricks is the same problem task as mentioned in the cafeteria scene. Not all models excelled at the same reasoning tasks. Furthermore, for the evaluation of Claude Sonnet 4, we do not use extended thinking, because this disables the necessary tool use.
(a) General scene synthesis comparison.
(b) Object placement comparison.
Figure 11: Mabyduck interface used for human evaluation.